OpenStudy (anonymous):

An investor invested a total of $1,300 in two mutal funds. One fund earns a 9% profit while the other earned a 4% profit. If the investor's total profit was $92, how much was invested in each mutal fund?

OpenStudy (mathstudent55):

Let x = amount invested at 9% Let y = amount invested at 4% x + y = 1300 0.09x + 0.04y = 92 Solve the system of equations

OpenStudy (anonymous):

I dont know what to do after that i am confused after this point

OpenStudy (mathstudent55):

Solve the first equation for x, x = 1300 - y Now insert what x is in for x in other equation: 0.09x + 0.04y = 92 0.09(1300 - y) +0.04y = 92 Distribute 0.09: 117 - 0.09y + 0.04y = 92 -0.05y = -25 y = 500 x + y = 1300 x + 500 = 1300 x = 800 Answer: $800 invested at 9% and $500 invested at 4%

OpenStudy (anonymous):

thank you i sort of get it now.

OpenStudy (mathstudent55):

yw